OpenStudy (anonymous):

Find the vertex h(t) = 1/3t^2 - 2t + 8

Nnesha (nnesha):

\[\huge\rm \frac{ -b }{ 2a }\] formula to find vertex

OpenStudy (anonymous):

I did that and got 3.03,4.97 which is wrong

Nnesha (nnesha):

A is what b is what ?

OpenStudy (anonymous):

a is 1/3 and b is -2

OpenStudy (anonymous):

which becomes positive 2/2(1/3)

Nnesha (nnesha):

okay right so how would you solve denominator ?

Nnesha (nnesha):

2 times 1/3= ?

OpenStudy (anonymous):

2/1 * 1/3 = 2/3 or 0.67

OpenStudy (anonymous):

then I know I'm supposed to plug it into the equation I just for some reason keep getting the answer wrong

Nnesha (nnesha):

\[\huge\rm \frac{ 2 }{ \frac{ 2 }{ 3 } }\] change division to multiplication by flipping fraction which is at the denominator for example \[\huge\rm \frac{ \frac{ x }{ y} }{\frac{ a }{ b } } = \frac{ x }{ y } \times \frac{ b }{ a }\]

Nnesha (nnesha):

does hat make sense ?

Nnesha (nnesha):

that*